A passenger in a car wants to check that the car’s speedometer is working correctly.
there are lamp posts on the side of the road.
The passenger knows that the lamp posts are 80 m apart.
Explain how the passenger can check that the speedometer is working correctly.
1
Expert's answer
2020-10-02T07:14:39-0400

He can record time that car takes to pass between two neighboring lamp post. Let's denote this time with "t". As we know, the distance between two neighboring lamp post is "h = 80m". Thus, in order to find car's speed, we should calculate the following expression:


"v = \dfrac{h}{t} = \dfrac{80}{t}\dfrac{m}{s}"

In this way the passenger can verify the speedometer.